# Equivalence Relations on Vertex Operator Algebras, II: Witt Equivalence and Orbifolds

### v1

@@ 2.3. Vertex Algebras
For any $c\in\mathbb{C}$, there is the universal Virasoro vertex operator algebra $V_{\text{Vir}}(c,0)$ of central charge $c$. Precisely when $c=c_{p,q}\coloneqq 1-6(p-q)^{2}/(pq)$ for coprime $p,q\geq 2$, does $V_{\text{Vir}}(c,0)$ contain a maximal proper ideal and the corresponding simple quotient is the simple Virasoro vertex operator algebra (or Virasoro minimal model) $L_{\text{Vir}}(c_{p,q},0)$, which we denote by $L(c_{p,q})$ for short ([FZ92, Wan93], see also [LL04]).
@@ 2.4. Physics
We recall the physical meaning of Witt equivalence of modular tensor categories, 2.10. Remembering that a TQFT $(\mathcal{C},c)$ admits a topological boundary condition if and only if $c=0$ and $\mathcal{C}$ is of the form $Z(\mathcal{F})$ for some spherical fusion category $\mathcal{F}$, we learn that $\mathcal{C}$ and $\mathcal{D}$ are Witt equivalent if and only if the theory $(\mathcal{C}\boxtimes\overline{\mathcal{D}},0)$ obtained by stacking $(\mathcal{C},c)$ onto the orientation-reversal of $(\mathcal{D},c)$ has a topological boundary. Or, by "unfolding" along the boundary, we can reformulate the Witt equivalence of $\mathcal{C}$ and $\mathcal{D}$ as the existence of a topological interface which interpolates between the TQFTs $(\mathcal{C},c)$ and $(\mathcal{D},c)$. This is depicted in Figure 7.
@@
Recall that we may think of a rational chiral algebra as a gapless chiral boundary condition of a 3d TQFT. Many statements about chiral algebras become clearer when viewed from this three-dimensional perspective. For example, the fact that a condensable algebra $A$ of $\operatorname{Rep}(W)$ defines a conformal extension $W\subset V$ with $\operatorname{Rep}(V)\cong\operatorname{Rep}(W)/A$ can be seen by taking the topological interface that interpolates between $\operatorname{Rep}(W)$ and $\operatorname{Rep}(W)/A$ (obtained by condensing anyons in half of spacetime) and fusing it onto the gapless chiral boundary condition described by $W$ to obtain a new gapless chiral boundary condition of $\operatorname{Rep}(W)/A$. This is depicted in Figure 8.
@@ 3.1. Weak Witt Equivalence
We begin by considering the weaker variation of Witt equivalence. To motivate our definition, we rephrase Witt equivalence of lattices, 2.3, in vertex algebraic terms. Recall the definition of Witt equivalence of modular tensor categories, 2.10, and also the definition of lattice vertex algebras or chiral free boson theories (see [Bor86, FLM88]).
@@
We only consider the forward direction. The logic of the reverse direction is nearly identical. Because $L$ and $M$ are assumed to be Witt equivalent, $c(V_{L})=\operatorname{rk}(L)=\operatorname{rk}(M)=c(V_{M})$. On the other hand, by the discussion around equation (1), $L^{\prime}/L$ and $M^{\prime}/M$ having isometric subquotients means that $L$ and $M$ can be extended to lattices $\tilde{L}$ and $\tilde{M}$ that have isometric discriminant forms. Equivalently, $V_{L}$ and $V_{M}$ can be conformally extended to lattice vertex operator algebras $V_{\tilde{L}}$ and $V_{\tilde{M}}$, respectively, which have ribbon equivalent representation categories, $\operatorname{Rep}(V_{\tilde{L}})\cong\operatorname{Rep}(V_{\tilde{M}})$. This means that there are condensable algebras $A$ and $B$ of $\operatorname{Rep}(V_{L})$ and $\operatorname{Rep}(V_{M})$, respectively, such that $\operatorname{Rep}(V_{L})_{A}^{\mathrm{loc}}$ is ribbon equivalent to $\operatorname{Rep}(V_{M})_{B}^{\mathrm{loc}}$. By item (2) of 2.11, it follows that $\operatorname{Rep}(V_{L})$ and $\operatorname{Rep}(V_{M})$ are Witt equivalent as modular tensor categories. ∎
@@
Evidently, two strongly rational vertex operator algebras $V$ and $V^{\prime}$ in the same bulk genus, 2.14, are Witt equivalent.
@@
The following gives an alternative characterization of Witt equivalence of strongly rational vertex operator algebras, analogous to item (2) of 2.11, which applies to modular tensor categories.
@@
Suppose that we are given $W$ and $W^{\prime}$ satisfying the hypotheses of the proposition. Evidently, $c(V)=c(V^{\prime})$. Furthermore, $W$ must be obtained via a condensable algebra $A$ in $\operatorname{Rep}(V)$ for which $\operatorname{Rep}(V)_{A}^{\mathrm{loc}}\cong\operatorname{Rep}(W)$, and similarly for $W^{\prime}$ (see 2.16). By item (2) of 2.11, $\operatorname{Rep}(V)$ is Witt equivalent to $\operatorname{Rep}(V^{\prime})$ as a modular tensor category.
Conversely, if $V$ and $V^{\prime}$ are Witt equivalent (and positive), then by item (2) of 2.11, there exist condensable algebras $A$ and $A^{\prime}$ in $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$, respectively, such that $\operatorname{Rep}(V)_{A}^{\mathrm{loc}}\cong\operatorname{Rep}(V^{\prime})_{A^{\prime}}^{\mathrm{loc}}$ as modular tensor categories. On the other hand, a condensable algebra $A$ in $\operatorname{Rep}(V)$ defines a conformal extension of $V$ to a strongly rational vertex operator algebra $W$ with $\operatorname{Rep}(W)\cong\operatorname{Rep}(V)_{A}^{\mathrm{loc}}$; similarly, one obtains a strongly rational vertex operator algebra $W^{\prime}$ containing $V^{\prime}$ as a conformal subalgebra and having $\operatorname{Rep}(W^{\prime})\cong\operatorname{Rep}(V^{\prime})_{A^{\prime}}^{\mathrm{loc}}$ (see 2.18). Moreover, $c(W)=c(V)=c(V^{\prime})=c(W^{\prime})$. ∎
@@
We depict the above characterization of Witt equivalence in Figure 10 (cf. Figure 4). In the special case of $W\cong W^{\prime}=V^{\prime}$ we obtain (cf. 2.12, Lemma 7.12 in [SY24]):
@@
Recall from 2.13 that every Witt class $[\mathcal{C}]$ of modular tensor categories admits a unique anisotropic representative $\mathcal{C}_{\mathrm{an}}$. Moreover, every modular tensor category $\mathcal{D}$ that is Witt equivalent to $\mathcal{C}_{\mathrm{an}}$ admits a (maximal) condensable algebra $A$ satisfying $\mathcal{C}_{\mathrm{an}}\cong\mathcal{D}_{A}^{\mathrm{loc}}$. The vertex operator algebra translation of this is the following.
@@
By 3.10, one only needs to check that $V$ is Witt equivalent to $V^{\prime}$ and that $c(W)=c(W^{\prime})$ to decide whether ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ are Witt equivalent as RCFTs. In particular, a Witt class of RCFTs can be labeled by a triple $(c_{L},c_{R},[\mathcal{C}])$ consisting of the left- and right-moving central charges $c_{L}$ and $c_{R}$, and the Witt class $[\mathcal{C}]$ of the representation category of the left-moving (or right-moving) chiral algebra of any representative RCFT of the class.
@@ 3.2. Strong Witt Equivalence
Now, since $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$ are Witt equivalent modular tensor categories by assumption, so are $\operatorname{Rep}(V)\boxtimes\overline{\operatorname{Rep}(C)}$ and $\operatorname{Rep}(V^{\prime})\boxtimes\overline{\operatorname{Rep}(C)}$. The same can be said about their condensations by 2.12. So, $\operatorname{Rep}(V_{L})\cong\mathcal{C}(L^{\prime}/L)$ and $\operatorname{Rep}(V_{M})\cong\mathcal{C}(M^{\prime}/M)$ are Witt equivalent, which by 3.1 means that the associated lattices $L$ and $M$ are Witt equivalent. Hence, $V$ and $V^{\prime}$ are in the same strong Witt class. ∎
@@ 3.3. Examples
By 3.17, in order to describe the decomposition of $([\operatorname{Vect}],c)$ into strong Witt classes, we need to simply add the Heisenberg commutant as a further invariant. It would be interesting to study this in detail for $c=24$. As a first step in this direction, in [HM23] we describe the possible Heisenberg commutants of index-$2$ conformal subalgebras of the holomorphic $c=24$ vertex operator algebras. In fact, we only do this for roughly half of them, namely the ones that correspond to non-anomalous orbifolds of order $2$. For these, one finds that, modulo 2.20, there are exactly $18$ Heisenberg commutants up to isomorphism.
@@
We revisit the minimal models in the discrete series from 2.19, i.e. the positive, strongly rational vertex operator algebras $L(c_{m})$ of central charge $c_{m}=1-\frac{6}{(m+2)(m+3)}<1$ and with (pseudo-unitary) representation category $\mathcal{M}_{m}=\operatorname{Rep}(L(c_{m}))$ for $m\in\mathbb{Z}_{\geq 0}$. In the following, we describe the strongly rational, positive vertex operator algebras that reside in the Witt classes $([\mathcal{M}_{m}],c_{m})$, as well as how those vertex operator algebras split into strong Witt classes.
@@
Furthermore, the special module $L(c_{m},h^{(m)}_{m+1,1})$ is always a simple-current module of $L(c_{m})$, and $\smash{h^{(m)}_{m+1,1}}=m(m+1)/4$ is an integer whenever $m\equiv 0,3\pmod{4}$. Then, by 2.18, there is a unique vertex operator algebra structure on
@@
We come to the case $m=1$. In that case, the lattice vertex operator algebra $V_{A_{1}}\cong\mathsf{A}_{1,1}$ is also the simple affine vertex operator algebra at level $1$. After all, $A_{1}$ is the root lattice of $\mathfrak{sl}_{2}$. Any lift of the $(-\operatorname{id})$-involution on the lattice $A_{1}$ is an inner automorphism of the vertex operator algebra $V_{A_{1}}$. Indeed, $V_{A_{1}}$ is strongly generated by the standard generators $\{e,f,h\}$ of the weight-$1$ Lie algebra $\mathfrak{sl}_{2}$ and the outer automorphism group of $\mathfrak{sl}_{2}$ is trivial. Hence, the automorphism group of $V_{A_{1}}$ is given by the inner automorphism group of $\mathfrak{sl}_{2}$, which is $\mathrm{SL}(2,\mathbb{C})$ acting by conjugation.
Now, the finite subgroups of $\mathrm{SL}(2,\mathbb{C})$ or equivalently of $\mathrm{SU}(2)$, up to conjugation, famously correspond bijectively to the Dynkin diagrams of ADE-type (via the Kleinian or Du Val singularities). The fixed-point vertex operator subalgebras of $V_{A_{1}}$ under subgroups of $\operatorname{Aut}(V_{A_{1}})$ corresponding to Dynkin types $A$ (cyclic subgroups) and $D$ (dihedral groups) are of the form $\smash{V_{\sqrt{k}A_{1}}}$ and $\smash{V_{\sqrt{k}A_{1}}^{+}}$, respectively, for $k/m$ a perfect square, as already discussed for $m>1$. The remaining three subgroups $G(E_{n})<\operatorname{Aut}(V_{A_{1}})$ of type $E_{6}$, $E_{7}$ and $E_{8}$ are exceptional cases called the tetrahedral, octahedral and icosahedral group, respectively. We denote the corresponding fixed-point vertex operator subalgebras by $V(E_{n})\coloneqq\smash{V_{A_{1}}^{G(E_{n})}}$. Except for $G(E_{8})$, all these groups are solvable so that we know that the fixed-point vertex operator subalgebras are strongly rational by [Miy15, CM16], with the possible exception of $V(E_{8})$. Of course, one is inclined to believe that also $V(E_{8})$ is strongly rational. Moreover, these fixed-point vertex operator subalgebras are all positive (provided they are strongly rational), as their irreducible modules are contained in twisted modules for $V_{A_{1}}$ [DRX17], which have positive weight grading [DL96].
@@
On the other hand, we can show that the lattice vertex operator algebras $V_{\sqrt{k}A_{1}}$ all belong to different Witt classes as one varies over squarefree integers $k\in\mathbb{Z}_{>0}$. To see this, recall that two inequivalent modular tensor categories belong to different Witt classes if they are both anisotropic since there is only one anisotropic modular tensor category in every Witt class, up to ribbon equivalence (see 2.13). Thus, since the categories $\smash{\operatorname{Rep}(V_{\sqrt{k}A_{1}})}\cong\mathcal{C}((\sqrt{k}A_{1})^{\prime}/(\sqrt{k}A_{1}))$ are ribbon inequivalent for different values of squarefree $k$ (they have different ranks, for example), it will follow that the vertex operator algebras $\smash{V_{\sqrt{k}A_{1}}}$ belong to different Witt classes if we prove that each category $\operatorname{Rep}(V_{\sqrt{k}A_{1}})$ is anisotropic.

### v2

@@ 2.3. Vertex Algebras
For any $c\in\mathbb{C}$, there is the universal Virasoro vertex operator algebra $V_{\mathrm{Vir}}(c,0)$ of central charge $c$. Precisely when $c=c_{p,q}\coloneqq 1-6(p-q)^{2}/(pq)$ for coprime $p,q\geq 2$, does $V_{\mathrm{Vir}}(c,0)$ contain a maximal proper ideal and the corresponding simple quotient is the simple Virasoro vertex operator algebra (or Virasoro minimal model) $L_{\mathrm{Vir}}(c_{p,q},0)$, which we denote by $L(c_{p,q})$ for short ([FZ92, Wan93], see also [LL04]).
@@ 2.4. Physics
We recall the physical meaning of Witt equivalence of modular tensor categories, 2.11. Remembering that a TQFT $(\mathcal{C},c)$ admits a topological boundary condition if and only if $c=0$ and $\mathcal{C}$ is of the form $Z(\mathcal{F})$ for some spherical fusion category $\mathcal{F}$, we learn that $\mathcal{C}$ and $\mathcal{D}$ are Witt equivalent if and only if the theory $(\mathcal{C}\boxtimes\overline{\mathcal{D}},0)$ obtained by stacking $(\mathcal{C},c)$ onto the orientation-reversal of $(\mathcal{D},c)$ has a topological boundary. Or, by "unfolding" along the boundary, we can reformulate the Witt equivalence of $\mathcal{C}$ and $\mathcal{D}$ as the existence of a topological interface which interpolates between the TQFTs $(\mathcal{C},c)$ and $(\mathcal{D},c)$. This is depicted in Figure 8.
@@
Recall that we may think of a rational chiral algebra as a gapless chiral boundary condition of a 3d TQFT. Many statements about chiral algebras become clearer when viewed from this $3$-dimensional perspective. For example, the fact that a condensable algebra $A$ of $\operatorname{Rep}(W)$ defines a conformal extension $W\subset V$ with $\operatorname{Rep}(V)\cong\operatorname{Rep}(W)/A$ can be seen by taking the topological interface that interpolates between $\operatorname{Rep}(W)$ and $\operatorname{Rep}(W)/A$ (obtained by condensing anyons in half of spacetime) and fusing it onto the gapless chiral boundary condition described by $W$ to obtain a new gapless chiral boundary condition of $\operatorname{Rep}(W)/A$. This is depicted in Figure 9.
@@ 3.1. Weak Witt Equivalence
We begin by considering the weaker variation of Witt equivalence. To motivate our definition, we rephrase Witt equivalence of lattices, 2.3, in vertex algebraic terms. Recall the definition of Witt equivalence of modular tensor categories, 2.11, and also the definition of lattice vertex algebras or chiral free boson theories (see [Bor86, FLM88]).
@@
We only consider the forward direction. The logic of the reverse direction is nearly identical. Because $L$ and $M$ are assumed to be Witt equivalent, $c(V_{L})=\operatorname{rk}(L)=\operatorname{rk}(M)=c(V_{M})$. On the other hand, by the discussion around equation (1), $L^{\prime}/L$ and $M^{\prime}/M$ having isometric subquotients means that $L$ and $M$ can be extended to lattices $\tilde{L}$ and $\tilde{M}$ that have isometric discriminant forms. Equivalently, $V_{L}$ and $V_{M}$ can be conformally extended to lattice vertex operator algebras $V_{\tilde{L}}$ and $V_{\tilde{M}}$, respectively, which have ribbon equivalent representation categories, $\operatorname{Rep}(V_{\tilde{L}})\cong\operatorname{Rep}(V_{\tilde{M}})$. This means that there are condensable algebras $A$ and $B$ of $\operatorname{Rep}(V_{L})$ and $\operatorname{Rep}(V_{M})$, respectively, such that $\operatorname{Rep}(V_{L})_{A}^{\mathrm{loc}}$ is ribbon equivalent to $\operatorname{Rep}(V_{M})_{B}^{\mathrm{loc}}$. By item (2) of 2.12, it follows that $\operatorname{Rep}(V_{L})$ and $\operatorname{Rep}(V_{M})$ are Witt equivalent as modular tensor categories. ∎
@@
Evidently, two strongly rational vertex operator algebras $V$ and $V^{\prime}$ in the same bulk genus, 2.15, are Witt equivalent.
@@
The following gives an alternative characterization of Witt equivalence of strongly rational vertex operator algebras, analogous to item (2) of 2.12, which applies to modular tensor categories.
@@
Suppose that we are given $W$ and $W^{\prime}$ satisfying the hypotheses of the proposition. Evidently, $c(V)=c(V^{\prime})$. Furthermore, $W$ must be obtained via a condensable algebra $A$ in $\operatorname{Rep}(V)$ for which $\operatorname{Rep}(V)_{A}^{\mathrm{loc}}\cong\operatorname{Rep}(W)$, and similarly for $W^{\prime}$ (see 2.17). By item (2) of 2.12, $\operatorname{Rep}(V)$ is Witt equivalent to $\operatorname{Rep}(V^{\prime})$ as a modular tensor category.
Conversely, if $V$ and $V^{\prime}$ are Witt equivalent (and positive), then by item (2) of 2.12, there exist condensable algebras $A$ and $A^{\prime}$ in $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$, respectively, such that $\operatorname{Rep}(V)_{A}^{\mathrm{loc}}\cong\operatorname{Rep}(V^{\prime})_{A^{\prime}}^{\mathrm{loc}}$ as modular tensor categories. On the other hand, a condensable algebra $A$ in $\operatorname{Rep}(V)$ defines a conformal extension of $V$ to a strongly rational vertex operator algebra $W$ with $\operatorname{Rep}(W)\cong\operatorname{Rep}(V)_{A}^{\mathrm{loc}}$; similarly, one obtains a strongly rational vertex operator algebra $W^{\prime}$ containing $V^{\prime}$ as a conformal subalgebra and having $\operatorname{Rep}(W^{\prime})\cong\operatorname{Rep}(V^{\prime})_{A^{\prime}}^{\mathrm{loc}}$ (see 2.19). Moreover, $c(W)=c(V)=c(V^{\prime})=c(W^{\prime})$. ∎
@@
We depict the above characterization of Witt equivalence in Figure 11 (cf. Figure 4). In the special case of $W\cong W^{\prime}=V^{\prime}$ we obtain (cf. 2.13, Lemma 7.12 in [SY24]):
@@
Recall from 2.14 that every Witt class $[\mathcal{C}]$ of modular tensor categories admits a unique anisotropic representative $\mathcal{C}_{\mathrm{an}}$. Moreover, every modular tensor category $\mathcal{D}$ that is Witt equivalent to $\mathcal{C}_{\mathrm{an}}$ admits a (maximal) condensable algebra $A$ satisfying $\mathcal{C}_{\mathrm{an}}\cong\mathcal{D}_{A}^{\mathrm{loc}}$. The vertex operator algebra translation of this is the following.
@@
By 3.10, one only needs to check that $V$ is Witt equivalent to $V^{\prime}$ and that $c(W)=c(W^{\prime})$ to decide whether ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ are Witt equivalent as RCFTs. In particular, a Witt class of RCFTs can be labeled by a triple $([\mathcal{C}],c_{L},c_{R})$ consisting of the left- and right-moving central charges $c_{L}$ and $c_{R}$, and the Witt class $[\mathcal{C}]$ of the representation category of the left-moving (or right-moving) chiral algebra of any representative RCFT of the class.
@@ 3.2. Strong Witt Equivalence
Now, since $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$ are Witt equivalent modular tensor categories by assumption, so are $\operatorname{Rep}(V)\boxtimes\overline{\operatorname{Rep}(C)}$ and $\operatorname{Rep}(V^{\prime})\boxtimes\overline{\operatorname{Rep}(C)}$. The same can be said about their condensations by 2.13. So, $\operatorname{Rep}(V_{L})\cong\mathcal{C}(L^{\prime}/L)$ and $\operatorname{Rep}(V_{M})\cong\mathcal{C}(M^{\prime}/M)$ are Witt equivalent, which by 3.1 means that the associated lattices $L$ and $M$ are Witt equivalent. Hence, $V$ and $V^{\prime}$ are in the same strong Witt class. ∎
@@ 3.3. Examples
By 3.17, in order to describe the decomposition of $([\operatorname{Vect}],c)$ into strong Witt classes, we need to simply add the Heisenberg commutant as a further invariant. It would be interesting to study this in detail for $c=24$. As a first step in this direction, in [HM23] we describe the possible Heisenberg commutants of index-$2$ conformal subalgebras of the holomorphic $c=24$ vertex operator algebras. In fact, we only do this for roughly half of them, namely the ones that correspond to non-anomalous orbifolds of order $2$. For these, one finds that, modulo 2.21, there are exactly $18$ Heisenberg commutants up to isomorphism.
@@
We revisit the minimal models in the discrete series from 2.20, i.e. the positive, strongly rational vertex operator algebras $L(c_{m})$ of central charge $c_{m}=1-6/((m+2)(m+3))<1$ and with (pseudo-unitary) representation category $\mathcal{M}_{m}=\operatorname{Rep}(L(c_{m}))$ for $m\in\mathbb{Z}_{\geq 0}$. In the following, we describe the strongly rational, positive vertex operator algebras that reside in the Witt classes $([\mathcal{M}_{m}],c_{m})$, as well as how those vertex operator algebras split into strong Witt classes.
@@
Furthermore, the special module $L(c_{m},h^{(m)}_{m+1,1})$ is always a simple-current module of $L(c_{m})$, and $\smash{h^{(m)}_{m+1,1}}=m(m+1)/4$ is an integer whenever $m\equiv 0,3\pmod{4}$. Then, by 2.19, there is a unique vertex operator algebra structure on
@@
We come to the case $m=1$. In that case, the lattice vertex operator algebra $V_{A_{1}}\cong\mathsf{A}_{1,1}$ is also the simple affine vertex operator algebra at level $1$. After all, $A_{1}$ is the root lattice of $\mathfrak{sl}_{2}$. Any lift of the $(-\operatorname{id})$-involution on the lattice $A_{1}$ is an inner automorphism of the vertex operator algebra $V_{A_{1}}$. Indeed, $V_{A_{1}}$ is strongly generated by the standard generators $\{e,f,h\}$ of the weight-$1$ Lie algebra $\mathfrak{sl}_{2}$ and the outer automorphism group of $\mathfrak{sl}_{2}$ is trivial. Hence, the automorphism group of $V_{A_{1}}$ is given by the inner automorphism group of $\mathfrak{sl}_{2}$, which is $\mathrm{PSL}(2,\mathbb{C})$ acting by conjugation.
Now, the finite subgroups of $\mathrm{PSL}(2,\mathbb{C})=\mathrm{SL}(2,\mathbb{C})/\{\pm\operatorname{id}\}$ or equivalently of $\mathrm{SU}(2)/\{\pm\operatorname{id}\}\cong\mathrm{SO}(3)$, up to conjugation, famously correspond bijectively to the Dynkin diagrams of ADE-type (via the Kleinian or Du Val singularities). The fixed-point vertex operator subalgebras of $V_{A_{1}}$ under subgroups of $\operatorname{Aut}(V_{A_{1}})$ corresponding to Dynkin types $A$ (cyclic subgroups) and $D$ (dihedral groups) are of the form $\smash{V_{\sqrt{k}A_{1}}}$ and $\smash{V_{\sqrt{k}A_{1}}^{+}}$, respectively, for $k$ a perfect square, as already discussed for $m>1$. The remaining three subgroups $G(E_{n})<\operatorname{Aut}(V_{A_{1}})$ of type $E_{6}$, $E_{7}$ and $E_{8}$ are exceptional cases called the tetrahedral, octahedral and icosahedral group, absractly isomorphic to $A_{4}$, $S_{4}$ and $A_{5}$, respectively. We denote the corresponding fixed-point vertex operator subalgebras by $V(E_{n})\coloneqq\smash{V_{A_{1}}^{G(E_{n})}}$. Except for $G(E_{8})\cong A_{5}$, all these groups are solvable so that we know that the fixed-point vertex operator subalgebras are strongly rational by [Miy15, CM16], with the possible exception of $V(E_{8})$. Of course, one is inclined to believe that also $V(E_{8})$ is strongly rational. Moreover, these fixed-point vertex operator subalgebras are all positive (provided they are strongly rational), as their irreducible modules are contained in twisted modules for $V_{A_{1}}$ [DRX17], which have positive weight grading [DL96].
@@
On the other hand, we can show that the lattice vertex operator algebras $V_{\sqrt{k}A_{1}}$ all belong to different Witt classes as one varies over squarefree integers $k\in\mathbb{Z}_{>0}$. To see this, recall that two inequivalent modular tensor categories belong to different Witt classes if they are both anisotropic since there is only one anisotropic modular tensor category in every Witt class, up to ribbon equivalence (see 2.14). Thus, since the categories $\smash{\operatorname{Rep}(V_{\sqrt{k}A_{1}})}\cong\mathcal{C}((\sqrt{k}A_{1})^{\prime}/(\sqrt{k}A_{1}))$ are ribbon inequivalent for different values of squarefree $k$ (they have different ranks, for example), it will follow that the vertex operator algebras $\smash{V_{\sqrt{k}A_{1}}}$ belong to different Witt classes if we prove that each category $\operatorname{Rep}(V_{\sqrt{k}A_{1}})$ is anisotropic.